OpenStudy (waheguru):

3(x-5)+4(3x+1)=2(6x-1)+11 evalutte

OpenStudy (waheguru):

the book says x= 13??????

OpenStudy (saifoo.khan):

Burn your book, the answer is not 13.

OpenStudy (anonymous):

Its 20/3

OpenStudy (waheguru):

wat is teh rea; snawerr

OpenStudy (saifoo.khan):

x = 20/3

OpenStudy (anonymous):

3(x-5)+4(3x+1)=2(6x-1)+11 3x -15 + 12x + 4 = 12x - 2 + 11 15x - 11 = 12x + 9 15x - 12x = 9 + 11 3x = 20 x = 20/3

OpenStudy (waheguru):

yea thats what i am getting so the book is wrong???

OpenStudy (anonymous):

i got x=8

OpenStudy (anonymous):

sometimes there are printing errors.....